use synthetic division and the given factor to completely factor each polynomial function:
y=x^3-4x^2-9x+36; (x+3)
HELP!!

Answer by lwsshak3(11628)

use synthetic division and the given factor to completely factor each polynomial function:
y=x^3-4x^2-9x+36; (x+3)
**
...-3...|...1.....-4......-9......36.........
..........|..........-3......21....-36
..........|...1.....-7......12......0..
y=(x+3)(x^2-7x+12)
y=(x+3)(x-4)(x-3)